OpenStudy (anonymous):

What is the value of x in the following equation: 3x + 21 - x = 2x + 20? 21 20 0 none of these

OpenStudy (amoodarya):

3x + 21 - x = 2x + 20 first simplify then put variable in left side and number on the other side

OpenStudy (awesome781):

What I do is plug in the values for and if they equal each other, then that is the correct answer

OpenStudy (haseeb96):

3x+21-x=2x+20 21+2x=2x+20 it has no solution So D is the correct answer

OpenStudy (anonymous):

@Haseeb96 Thanks :)